Parramore Corp has $14 million of sales, $1 million of inventories, $4 million of receivables, and $1 million of payables. Its cost of goods sold is 70% of sales, and it finances working capital with bank loans at an 6% rate. Assume 365 days in year for your calculations. Do not round intermediate steps.

  1. What is Parramore's cash conversion cycle (CCC)? Do not round intermediate calculations. Round your answer to two decimal places.
      days

  2. If Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold, what would be the new CCC? Do not round intermediate calculations. Round your answer to two decimal places.
      days

  3. How much cash would be freed up, if Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
    $

  4. By how much would pretax profits change, if Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
Answer #1

Part a

days outstanding sale = 4*365 / 14 = 104.2857

days inventory outstanding = 1*365/ 9.8 = 37.24

days payable outstanding = 1*365/ 9.8 = 37.24

CASH Conversion cycle = 104.2857 + 37.24 -37.24 = 104.2857 or 104.29

Part b

days outstanding sale = 3.6*365 / 14 = 93.86

days inventory outstanding = 0.9*365/ 9.8 = 33.52

days payable outstanding = 1.1*365/ 9.8 = 40.97

CASH Conversion cycle = 93.86 + 33.52 -40.97= 86.41

Part c

Cash freed up = [4-3.6] + [1-0.9] + [1.1-1] = 0.6 = 600,000

Part d

Increase in pretax profit = 600,000 * 0.06 = 36,000